40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 21 18:40:46 2014, manasa, Update, General, End laser fiber setup 
    Reply  Fri Oct 24 10:08:24 2014, manasa, Update, General, Y AUX laser - fiber coupled 
       Reply  Fri Oct 24 18:53:23 2014, rana, Update, General, Y AUX laser - fiber coupled 
       Reply  Mon Oct 27 12:17:46 2014, manasa, Update, General, Y AUX laser coupling telescope design AUXmode.zipAUXmode.png
          Reply  Wed Oct 29 11:45:11 2014, manasa, Update, General, Y AUX laser - fiber coupled (52%) 31.png
Message ID: 10639     Entry time: Fri Oct 24 18:53:23 2014     In reply to: 10638
Author: rana 
Type: Update 
Category: General 
Subject: Y AUX laser - fiber coupled 

 10% seems like a pretty bad coupling efficiency, even for a single lens. I know that the NPRO itself isn't so elliptical as that. Where is the other 230 mW going? random scattering?

Given that this is such an invasive process and, since its so painful to lose a whole night of locking due to end table business, I suggest that you always measure the out-of-loop ALS noise at the end of the end table work. Just checking that the green laser is locked to the arm is not sufficient to prove that the end table work won't prevent us from locking the interferometer.

We should insist on this anytime someone works on the optics or electronics at EX or EY. Don't have enough time to do an out-of-loop ALS spectrum? Then don't work at the end tables at all that day. We've got PZT alignment and mode matching work to do, as well as the rebuild of the EX table enclosure, so this is a good discipline to pick up now.

ELOG V3.1.3-